2. Write the slope intercept from the equation that passes through the points (2,7) and (-1,-11).
Answer by texttutoring(324)

First use the equation for the slope of a line:

slope = m = (y2-y1 / x2-x1)

m = (-11 - 7 / -1 - 2)
m = -18/-3
m=6

Now use the equation of a line, y=mx+b. You now know m, and you know y and x from either of the two points (2,7) or (-1,-11)

y=mx+b
7=6*2 + b
7=12+b
7-12=b
-5 = b

Putting it all together, with m=6 and b=-5, the equation of the line is:

y=6x-5